# Spin-operator form factors of the critical Ising chain and their finite volume scaling limits

### v3

@@ Acknowledgements.
The author thanks Zoltán Bajnok for constructive discussions and introduction to the references Yurov and Zamolodchikov (1991); Fonseca and Zamolodchikov (2001).
Appendix A The $s\rightarrow 0^{+}$ expansion of the crossover function
In this appendix, we compute analytically the small $s\rightarrow 0^{+}$ expansion of the function $g_{I}(s)$ of Sachdev (1996). We start from the definition with $s>0$
@@ Appendix A The $s\rightarrow 0^{+}$ expansion of the crossover function
In particular, the leading term exactly reproduces the constant ${\cal C}_{1}$ in Sachdev (1996) extracted numerically. At higher orders, there will be $\zeta_{2k+1}\zeta_{2l+1}$ type terms. The expansion has a convergence radius equals to $\pi$.

### v4

@@ Acknowledgements.
The author thanks Zoltán Bajnok for constructive discussions and introduction to the references Yurov and Zamolodchikov (1991); Fonseca and Zamolodchikov (2001). In particular, the author was alerted that there seems to be no explicit formulas for the CFT matrix elements in Yurov and Zamolodchikov (1991) after discussing with him. The results Eq. (109), Eq. (110) and Eq. (111) for the CFT matrix-elements were double checked by two less rigorous semi-infinite versions of this work, corresponding to taking the scaling limit at two different early stages. Notes for these computations can be provided upon request. The product formulas were checked by Zoltán Bajnok against the algorithm of Yurov and Zamolodchikov (1991) with his private codes up to phase conventions.
Appendix A The $s\rightarrow 0^{+}$ expansion of a crossover function
In this appendix, we compute analytically the small $s\rightarrow 0^{+}$ expansion of the function $g_{I}(s)$ of Sachdev (1996) as an application of the Barnes integral Eq. (219). We start from the definition with $s>0$
@@ Appendix A The $s\rightarrow 0^{+}$ expansion of a crossover function
In particular, the leading term exactly reproduces the constant ${\cal C}_{1}$ extracted numerically in Sachdev (1996). It is also in agreement with the Eq. (5.19) of Bugrij (2001). At higher orders, there will be $\zeta_{2k+1}\zeta_{2l+1}$ type terms. The expansion has a convergence radius equals to $\pi$.